### v2

# On Difference Sets of Dense Subsets of $\mathbb{Z}^{2}$

Abstract
In this article, we study the structure of the difference set EEE-E for subsets $E\subseteq\mathbb{Z}^{2}$ of positive upper Banach density. Fish asked in [5, Problem 2] whether, for every such set $E$, there exists a nonzero integer $k$ such that
$k\cdot\mathbb{Z}\subseteq\{\,xy:(x,y)\in E-E\,\}.$
Although this question remains open, we establish a relatively weaker form of this conjecture. Specifically, we prove that if $\langle a_{j}\rangle_{j=1}^{m}$ is any finite sequence in $\mathbb{N},$ then there exist infinitely many integers $k\in\mathbb{Z}$ and a sequence $\langle x_{n}\rangle_{n\in\mathbb{N}}$ in $\mathbb{Z}$ such that
$k\cdot MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq\{\,xy:(x,y)\in E-E\,\},$
where $MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)$ denotes the milliken-Taylor configuration generated by the sequences $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n\in\mathbb{N}}$.
Mathematics subject classification 2020: 05D10, 11E25, 11T30.
Keywords: Sum-product estimates, Difference set, Product sets.
1  Introduction
Estimating the sizes of sumsets and product sets of subsets of $\mathbb{Z}$ is a central topic in additive combinatorics. For any two sets $A,B\subset\mathbb{Z}$ (or $\mathbb{R}$), we define their sumset and product set by
$A+B=\{a+b:a\in A,\ b\in B\}\quad\text{and}\quad A\cdot B=\{ab:a\in A,\ b\in B\}.$
One of the most fundamental open problems in sum–product theory is the following conjecture of Erdős and Szemerédi [4].
Conjecture 1.1 (Erdős–Szemerédi [4]).
If $A\subset\mathbb{Z}$ (or $\mathbb{R}$) is a finite set, then for every $\epsilon>0$,
$|A+A|+|A\cdot A|\gg|A|^{2-\epsilon}.$
However, if one replaces the finite set $A$ by an infinite one, then an immediate question appears what kind of structures are preserved under the combination of these sum-product of sets. For various questions on sum, difference, and product sets, we refer to the article [9]. In [3, 5], Bjőrklund and Fish initiated an infinitary version of these types of questions. For some recent developments, we refer to the article [2].
Definition 1.2 (Upper Banach density in $\mathbb{Z}^{2}$).
Let $A\subseteq\mathbb{Z}^{2}$. The upper Banach density of $A$ is defined by
$d^{\star}(A):=\lim_{N\to\infty}\sup_{(m,n)\in\mathbb{Z}^{2}}\frac{\left|A\cap\bigl((m,n)+[-N,N]^{2}\bigr)\right|}{(2N+1)^{2}}.$
In [5], Fish asked the following question.
Problem 1.3.
[5, Problem 2 (page 5)] Is it true that for every set $E\subset\mathbb{Z}^{2}$ of positive upper Banach density, there exists a constant $k_{0}$, depending only on $d^{*}(E)$, such that for some integer $k\leq k_{0}$,
$k\mathbb{Z}\subseteq\{\,xy:(x,y)\in E-E\,\}?$
Although this problem is still open, we prove a relatively weaker version of this result. But before we state our result, let us recall the notions of $IP$ sets. Let $\mathcal{P}_{f}\left(\mathbb{N}\right)$ be the collection of nonempty finite subsets of $\mathbb{N}.$ In any commutative semigroup $(S,+),$ a set $A$ is said to be an $IP$ set if there exists a sequence $\langle x_{n}\rangle_{n}$ such that
$A=FS\left(\langle x_{n}\rangle_{n\in\mathbb{N}}\right)=\left\{\sum_{t\in H}x_{t}:H\in\mathcal{P}_{f}\left(\mathbb{N}\right)\right\}.$
Over the semigroup $(\mathbb{N},\cdot),$ we denote it by $FP\left(\langle x_{n}\rangle_{n\in\mathbb{N}}\right).$ The notions of $IP$ sets have significant importance in Ramsey theory, for ex. it is known that they are partition regular (see [6]).
An extension of Hindman’s Finite Sum Theorem was studied independently by Milliken and Taylor in [8, 10]. To address their results, we need the following notions of Milliken-Taylor system.
Definition 1.4 (Milliken-Taylor system).
Let $m\in\mathbb{N},$ let $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n}$ be sequences in $\mathbb{N}.$ The Milliken-Taylor System determined by $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n}$ is
$MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)=\left\{\sum_{j=1}^{m}a_{j}\cdot\sum_{t\in F_{j}}x_{t}:F_{1}<F_{2}<\cdots<F_{m},\forall F_{i}\in\mathcal{P}_{f}\left(\mathbb{N}\right)\right\}.$
So clearly, for any sequence $\langle x_{n}\rangle_{n},$ $FS(\langle x_{n}\rangle_{n})=MT\left(\{1\},\langle x_{n}\rangle_{n}\right).$ However, the following theorem says a more general result than the Finite Sum Theorem. Although in the original text the result was described over the set of positive integers, the technique for $\mathbb{Z}$ is similar.
Theorem 1.5.
[7, Theorem 17.31] Let $\langle a_{j}\rangle_{j=1}^{m}$ be sequences in $\mathbb{N}.$ Then for any $p\in E(\beta\mathbb{Z},+)$ and any $A\in a_{1}p+\cdots+a_{m}p,$ there exists a sequence $\langle x_{n}\rangle_{n}$ such that $MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq A.$
In this article, we prove the following theorem, which is a weaker version of the Problem 1.3.
Theorem 1.6.
Let $m\in\mathbb{N},$ and $\langle a_{j}\rangle_{j=1}^{m}$ be a sequence in $\mathbb{N}$ is given. If $E\subset\mathbb{Z}^{2}$ is a set of positive upper Banach density, then there exist infinitely many $k\in\mathbb{Z}$ and infinitely many sequences $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ such that
$k\cdot MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq\{\,xy:(x,y)\in E-E\,\}.$
For any sequence $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ and $H\in\mathcal{P}_{f}\left(\mathbb{N}\right),$ define $x_{H}=\sum_{t\in H}x_{t}.$ Our technique also proves the following result.
Theorem 1.7.
If $E\subset\mathbb{Z}^{2}$ is a set of positive upper Banach density, then there exist infinitely many $k\in\mathbb{Z}$ and infinitely many sequences $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ such that
$\{x_{K}-x_{H}:H,K\in\mathcal{P}_{f}\left(\mathbb{N}\right),\text{ and }\max H<\min K\}\subseteq\{\,x+y:(x,y)\in E-E\,\}.$
2  Preliminaries
Ergodic preliminaries
In this article, we intend to study these notions of recurrence for countable amenable groups. The reason for restricting to countable amenable groups is that they provide a natural framework for Furstenberg’s correspondence principle.
Definition 2.1.
A discrete group $G$ is said to be amenable if there exists an invariant mean on the space $\mathcal{B}(G)$ of real-valued bounded functions on $G$, that is, a positive linear functional $L:\mathcal{B}(G)\to\mathbb{R}$ satisfying:
  1. 1.
    $L(\mathbf{1}_{G})=1$;
  2. 2.
    $L(f_{g})=L({}_{g}f)=L(f)$ for all $f\in\mathcal{B}(G)$ and all $g\in G$, where $f_{g}(t)=f(tg)$ and $_{g}f(t)=f(gt)$.
The existence of an invariant mean is only one of many equivalent characterizations of amenability. In what follows, we shall make frequent use of the Følner characterization of amenability for discrete groups, which will be particularly useful for our purposes.
Theorem 2.2.
A countable group $G$ is amenable if and only if it admits a left Følner sequence, that is, a sequence of finite nonempty sets $(F_{n})_{n\in\mathbb{N}}\subseteq G$ such that for every $g\in G$,
$\lim_{n\to\infty}\frac{|gF_{n}\triangle F_{n}|}{|F_{n}|}=0.$
Definition 2.3.
Let $G$ be a countable amenable group. A subset $E\subseteq G$ is said to have
  1. 1.
    positive Følner density, or positive upper density with respect to a left Følner sequence $(F_{n})_{n\in\mathbb{N}}$, if
    $\limsup_{n\to\infty}\frac{|E\cap F_{n}|}{|F_{n}|}>0.$
    This quantity will be denoted by $d_{F_{n}}(E)$.
  2. 2.
    The upper Banach density of $E$ is defined as $d^{\star}(E)=\sup_{(F_{n})}d_{F_{n}}(E),$ where the supremum is taken over all Følner sequences $(F_{n})$ in $G$.
Note that Central sets can be defined as members of minimal idempotent ultrafilters, and Central* sets are those sets that sit in every minimal idempotent ultrafilter. We will recall these notions in the coming section The following theorem, which follows from [1, Theorem 1.4], shows that the set of return times is large.
Theorem 2.4.
Let $G$ be a countable amenable group, let $\{T_{g}\}_{g\in G}$ and $\{S_{g}\}_{g\in G}$ be commuting measure preserving actions of $G$ on a probability space $(X,\mathcal{B},\mu)$, and let $A\in\mathcal{B}$ satisfy $\mu(A)>0$. Then the set
$\{g\in G:\mu(A\cap T_{g}^{-1}A\cap(T_{g}S_{g})^{-1}A)>0\}$
is Central* in $G$.
From the above theorem it is also clear that $\{g\in G:\mu(A\cap S_{g}^{-1}A\cap(T_{g}S_{g})^{-1}A)>0\}$ is Central* in $G$. Now, from the Furstenberg correspondence principle over amenable groups, and Theorem 2.4, it follows the following corollary.
Corollary 2.5.
Let $G$ be a countable left amenable group, and $E\subseteq G\times G$ be such that $d^{\star}(E)>0.$ Then
$\{x\in G:d\bigl(E\cap(x,e)^{-1}E\cap(x,x)^{-1}E\bigr)>0\}$
is a central* set. And similarly,
$\{x\in G:d\bigl(E\cap(e,x)^{-1}E\cap(x,x)^{-1}E\bigr)>0\}$
is a central* set.
Here, we define the notions of density recurrent sets, which will be useful in our study.
Definition 2.6.
Let $G$ be a countable left amenable group.
  1. 1.
    Let $B\subseteq G$. Then $B$ is left density $2$-recurrent if and only if whenever $E\subseteq G\times G$ satisfies $d(E)>0$, there exists $x\in B$ such that
    $d\bigl(E\cap(x,e)^{-1}E\cap(x,x)^{-1}E\bigr)>0.$
  2. 2.
    Let $C\subseteq G$. Then $B$ is right density $2$-recurrent if and only if whenever $E\subseteq G\times G$ satisfies $d(E)>0$, there exists $x\in B$ such that
    $d\bigl(E\cap(e,x)^{-1}E\cap(x,x)^{-1}E\bigr)>0.$
  3. 3.
    Let $D\subseteq G$. Then $B$ is both sided density $2$-recurrent if and only if whenever $E\subseteq G\times G$ satisfies $d(E)>0$, there exists $x\in B$ such that
    1. (a)
      $d\bigl(E\cap(e,x)^{-1}E\cap(x,x)^{-1}E\bigr)>0;\text{ and }$
    2. (b)
      $d\bigl(E\cap(x,e)^{-1}E\cap(x,x)^{-1}E\bigr)>0.$
A brief review of Topological Algebra
Here we recall some basic facts about ultrafilters. For details, refer to the book [7].
Let $(S,\cdot)$ be any discrete semigroup. Let $\beta S$ be the collection of all ultrafilters. For every $A\subseteq S,$ define $\overline{A}=\{p\in\beta S:A\in p\}.$ Now one can check that the collection $\{\overline{A}:A\subseteq S\}$ forms a basis for a topology. This basis generates a topology over $\beta S.$ We can extend the operation $\cdot$" of $S$ over $\beta S$ as: for any $p,q\in\beta S,$ $A\in p\cdot q$ if and only if $\{x:x^{-1}A\in q\}\in p,$ where $x^{-1}A=\{y\in S:xy\in A\}.$ With this operation $\cdot$", $(\beta S,\cdot)$ becomes a compact Hausdorff right topological semigroup. One can show that $\beta S$ is nothing but the Stone-Čech compactification of $S.$ Hence Ellis’s theorem guarantees that there exist idempotents in $(\beta S,\cdot)$. The set of all idempotents in $(\beta S,\cdot)$ is denoted by $E\left((\beta S,\cdot)\right).$ It can be shown that every member of the idempotents of $(\beta S,\cdot)$ contains an $IP$ set, which means every idempotent witnesses Hindman’s theorem. Using Zorn’s lemma, one can show that $(\beta S,\cdot)$ contains minimal left ideals (minimal w.r.t. the inclusion). A well-known fact is that the union of such minimal left ideals is a minimal two-sided ideal, denoted by $K(\beta S,\cdot).$ If our $S$ is a group, then for any ultrafilter $p$ define $p^{-1}=\{A^{-1}:A\in p\}.$ It can be verified that $p^{-1}$ is also an ultrafilter. Here we recall a few well-known classes of sets that are relevant to our work.
Definition 2.7.
  1. 1.
    AA is central set if it belongs to a minimal idempotent in $\beta S$.
  2. 2.
    AA is central* if it belongs to all minimal idempotents in $\beta S$.
The tensor product of two ultrafilters is defined as follows.
Definition 2.8.
Let $(S,\cdot)$ and $(T,\cdot)$ be two discrete semigroups. Let $p\in(\beta S,\cdot)$ and $q\in(\beta T,\cdot)$ be two ultrafilters. Then the tensor product of $p$ and $q$ is defined as
$p\otimes q=\{A\subseteq S\times T:\{x\in S:\{y:(x,y)\in A\}\in q\}\in p\}$
where $x\in S$, $y\in T.$
3  Our Results
For any countable amenable group $(G,\cdot),$ define the following notions.
Definition 3.1.
Define
  1. 1.
    $\mathcal{LDR}_{2}(G)=\{p\in\beta G:(\forall B\in p)\;B\text{ is left density $2$-recurrent}\};$
  2. 2.
    $\mathcal{RDR}_{2}(G)=\{p\in\beta G:(\forall B\in p)\;B\text{ is right density $2$-recurrent}\};$
  3. 3.
    $\mathcal{DR}_{2}(G)=\{p\in\beta G:(\forall B\in p)\;B\text{ is both density $2$-recurrent}\}.$
To obtain Theorem 1.6 and Theorem 1.7, we will study the properties of the set $AA^{-1}$, where $A\subseteq G\times G$ with positive upper Banach density. First, we prove a sequence of lemmas.
Lemma 3.2.
Let $G$ be a countable amenable group. Then all $\mathcal{LDR}_{2}(G)$ and $\mathcal{RDR}_{2}(G)$ are subsemigroups of $\beta G$ containing the minimal idempotents of $\beta G$.
Proof.
Here we prove the result $\mathcal{LDR}_{2}(G).$ However the proof for $\mathcal{RDR}_{2}(G)$ is similar and we leave it. By 2.5, whenever $E\subseteq G\times G$ satisfies $d^{\star}(E)>0$, the set
$\{x\in G:d^{\star}(E\cap(x,e)^{-1}E\cap(x,x)^{-1}E)>0\}$
is central*.
Let $p$ be a minimal idempotent in $\beta G$ and let $C\in p$. Then for every $E\subseteq G\times G$ with $d^{\star}(E)>0$,
$C\cap\{x\in G:d^{\star}(E\cap(x,e)^{-1}E\cap(x,x)^{-1}E)>0\}\neq\varnothing.$
Hence, there exists $g\in C$ such that
$d^{\star}(E\cap(g,e)^{-1}E\cap(g,g)^{-1}E)>0.$
Thus $C$ is density $2$-recurrent, which implies $K(\beta G)\subseteq\mathcal{DR}_{2}(G)$. In particular, $\mathcal{DR}_{2}(G)$ is nonempty.
Now let $p,q\in\mathcal{DR}_{2}(G)$. Let $B\in pq$ and $A\subseteq G\times G$ with $d^{\star}(A)>0$. Define
$C=\{x\in G:x^{-1}B\in q\}.$
Then $C\in p$, and hence $C$ is density $2$-recurrent. Choose $x\in C$ such that
$d^{\star}(A\cap(x,e)^{-1}A\cap(x,x)^{-1}A)>0,$
and set
$D=A\cap(x,e)^{-1}A\cap(x,x)^{-1}A.$
Since $x^{-1}B\in q$, there exists $y\in x^{-1}B$ such that
$d^{\star}(D\cap(y,e)^{-1}D\cap(y,y)^{-1}D)>0.$
As $xy\in B$, it suffices to show that
$D\cap(y,e)^{-1}D\cap(y,y)^{-1}D\subseteq A\cap(xy,e)^{-1}A\cap(xy,xy)^{-1}A.$
Let $(u,v)$ belong to the left-hand side. Then $(u,v)\in D\subseteq A$ and
$(u,v)\in(y,e)^{-1}D\subseteq(y,e)^{-1}(x,e)^{-1}A,$
so $(xyu,v)\in A$, which implies $(u,v)\in(xy,e)^{-1}A$. Similarly,
$(u,v)\in(y,y)^{-1}D\subseteq(y,y)^{-1}(x,x)^{-1}A,$
so $(xyu,xyv)\in A$, and hence $(u,v)\in(xy,xy)^{-1}A$. Therefore,
$(u,v)\in A\cap(xy,e)^{-1}A\cap(xy,xy)^{-1}A.$
Note that from corollary 2.5, it is clear that $E(K(\beta G,\cdot))\subseteq\mathcal{DR}_{2}(G).$ Although we don’t know for $p\in\mathcal{LDR}_{2}(G)$ if $p^{-1}\in\mathcal{LDR}_{2}(G),$ or the same for $\mathcal{RDR}_{2}(G)$. But here we show this result holds for $\mathcal{DR}_{2}(G).$
Lemma 3.3.
If $p\in\mathcal{DR}_{2}(G),$ then $p^{-1}\in\mathcal{DR}_{2}(G).$
Proof.
Let $p\in\mathcal{DR}_{2}(G),$ and $B\in p.$ It will be sufficient to show that $B^{-1}$ is a both density $2$-recurrent set. Pick any set $A$ with $d(A)>0.$ As $p\in\mathcal{DR}_{2}(G),$ there exists $x\in B$ such that $d(A\cap(x,e)^{-1}A\cap(x,x)^{-1}A)>0$ and $d(A\cap(e,x)^{-1}A\cap(x,x)^{-1}A)>0.$ Define
  1. 1.
    $D_{1}=A\cap(x,e)^{-1}A\cap(x,x)^{-1}A;$
  2. 2.
    $D_{2}=A\cap(e,x)^{-1}A\cap(x,x)^{-1}A.$
Let
  1. 1.
    $E_{1}=(x,x)\cdot D_{1}=A\cap(e,x)\cdot A\cap(x,x)\cdot A;$ and
  2. 2.
    $E_{2}=(x,x)\cdot D_{2}=A\cap(x,e)\cdot A\cap(x,x)\cdot A.$
Then both $E_{1}$ and $E_{2}$ have positive density. Hence $d(A\cap(e,x^{-1})^{-1}A\cap(x^{-1},x^{-1})^{-1}A)>0$ and $d(A\cap(x^{-1},e)^{-1}A\cap(x^{-1},x^{-1})^{-1}A)>0.$ This completes the proof. ∎
As both $\mathcal{LDR}_{2}(G)$ and $\mathcal{RDR}_{2}(G)$ contains $\mathcal{DR}_{2}(G),$ from lemma 3.2 and lemma 3.3 it is clear that if $p\in\mathcal{DR}_{2}(G)$ and $q\in\mathcal{LDR}_{2}(G),$ then $p^{-1}q\in\mathcal{LDR}_{2}(G).$ The same conclusion holds for $\mathcal{RDR}_{2}(G).$
Let
$T=\bigcap\left\{\overline{EE^{-1}}:A\subseteq G\times G\,\text{ and }\,d^{\star}(E)>0\right\}.$
The following lemma shows that the tensor product of two elements from $\mathcal{DR}_{2}(G)$ is a member of $T.$
Lemma 3.4.
If $p,q\in\mathcal{DR}_{2}(G),$ then $p\otimes q\in T.$
Proof.
Here, we only prove $p\otimes q\in T$ rest proofs are similar. If possible, assume that $p\otimes q\notin T.$ Hence there exists a set $A$ with $d(A)>0$ such that $A^{-1}A\notin p\otimes q.$ Hence $B=(A^{-1}A)^{c}\in p\otimes q.$ But $B\in p\otimes q$ implies $C=\{x:\{y:(x,y)\in B\}\in q\}\in p.$ Choose $x\in C$ such that $d(A\cap(x,e)^{-1}A\cap(x,x)^{-1}A)>0.$ Let
$D=A\cap(x,e)^{-1}A\cap(x,x)^{-1}A.$
Now there exists $y\in B_{x}=\{(y:(x,y)\in B\}$ such that $d(D\cap(e,y)^{-1}D\cap(y,y)^{-1}D)>0.$ But this implies $d(A\cap(x,y)^{-1}A)>0.$ But this implies $(x,y)\in A^{-1}A\cap B,$ a contradiction. ∎
Lemma 3.5.
The above proof suggests something more: if $p\in\mathcal{LDR}_{2}(G)$ and $q\in\mathcal{RDR}_{2}(G),$ then $p\otimes q\in T.$
That means if $E\subseteq\mathbb{Z}\times\mathbb{Z}$ is a subset with positive upper Banach density, then for any $p\in\mathcal{LDR}_{2}(G)$ and $q\in\mathcal{RDR}_{2}(G),$ we have $(E-E)\in p\otimes q.$
To prove Theorem 1.6, we need the following lemma. Again, the reference contains the proof over the set of positive integers, but the same proof holds over the set of all integers.
Lemma 3.6.
[7, Lemma 5.19.2] If $p\in E(K(\beta\mathbb{Z},+))$ and $n\in\mathbb{N},$ then $n\cdot p\in E(K(\beta\mathbb{Z},+)).$
Now we prove our first Theorem.
Proof of Theorem 1.6.
Define the map $f:\mathbb{Z}^{2}\to\mathbb{Z}$ by $f(x,y)=xy.$ Let $\beta f:\beta\mathbb{Z}^{2}\to\beta\mathbb{Z}$ be its continuous extension. First, we claim that $\beta f(p\otimes q)=p\cdot q.$ As both are ultrafilters, it will be sufficient to show that $\beta f(p\otimes q)\subseteq p\cdot q.$ Suppose $B\in\beta f(p\otimes q).$ Then $f^{-1}(B)\in p\otimes q.$ Then $\{a:\{b:(a,b)\in f^{-1}(B)\}\in q\}\in p.$ But $(a,b)\in f^{-1}(B)\implies a\cdot b\in B.$ But this implies $B\in p\cdot q.$ And this proves the claim.
Now choose any $p,q\in E(K(\beta\mathbb{Z},+))$. Then from Lemmas 3.6, 3.2 we have $a_{1}q+\cdots+a_{m}q\in\mathcal{RDR}_{2}(\mathbb{Z}).$ But from Lemma 3.5 and our claim, it follows that
$f(E-E)\in\beta f(p\otimes(a_{1}q+\cdots+a_{m}q))=p\cdot(a_{1}q+\cdots+a_{m}q).$
Hence
$\{\,xy:(x,y)\in E-E\,\}\in p\cdot(a_{1}q+\cdots+a_{m}q).$
Now from Theorem 1.5, it is an routine exercise to check that sequences $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ such that
$k\cdot MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq\{\,xy:(x,y)\in E-E\,\}.$
Proof of Theorem 1.7.
Suppose $E\subseteq\mathbb{Z}^{2}$ is the set with positive upper Banach density. Define the map $f:\mathbb{Z}^{2}\to\mathbb{Z}$ by $f(x,y)=x+y.$ Let $\beta f:\beta\mathbb{Z}^{2}\to\beta\mathbb{Z}$ be its continuous extension. Likewise, the above proof it is immediate that for any $p,q\in\beta\mathbb{Z},$ $\beta f(p\otimes q)=p+q.$ Now choose any $p\in E(K(\beta\mathbb{Z},+)).$ Then from Lemmas 3.3, 3.5 it follows that $(E-E)\in(-p)\otimes p.$ And so,
$f(E-E)=\{\,x+y:(x,y)\in E-E\,\}\in(-p)+p.$
So to prove our result, it will be sufficient to show that all the members of (p)+p(-p)+p contain the desired pattern. Let $A\in(-p)+p.$ Then $\{x:-x+A\in p\}\in(-p).$ Hence $B=\{y:y+A\in p\}\in p.$ Note $B^{\star}=\{y:-y+B\in p\}\in p.$ Pick $y_{1}\in B^{\star}.$ From [7, Lemma 4.14], $y_{1}+B^{\star}\in p.$ Then pick
$y_{2}\in B_{2}=B^{\star}\cap(-y_{1}+B^{\star})\cap(y_{1}+A)\in p.$
Hence $y_{2}-y_{1}\in A.$ Now pick
$y_{3}\in B^{\star}\cap\bigcap_{y\in FS({y_{1},y_{2}})}(-y+B^{\star})\cap\bigcap_{y\in FS({y_{1},y_{2}})}(y+A)\cap(-y_{2}+B_{2}^{\star})\cap(-y_{1}+B_{1}^{\star})\in p.$
Then clearly for any $y\in FS({y_{1},y_{2}})$, $y_{3}-y\in A.$ Also from the choice $y_{3}\in-y_{2}+B_{2}^{\star},$ we have $y_{2}+y_{3}-y_{1}\in A.$ And continuing this iteration method, one can prove our result.
Acknowledgment
The author of this paper is supported by NBHM postdoctoral fellowship with reference no: 0204/27/(27)/2023/R & D-II/11927.